# Design and Implementation of a Scalable Clinical Data Warehouse for Resource-Constrained Healthcare Systems

### v1

@@ Abstract
Index Terms— dengue, datamart, health informatics, data mining, warehouse, electronic health records, infectious disease, outbreak detection
@@ I Introduction & Related Works
A National Clinical Data Warehouse (NCDW) offers a viable approach to consolidating multi-institutional data into a privacy-preserving environment. By unifying patient information, laboratory results, and clinical observations, an NCDW can enhance large-scale health analytics and support more informed policy decisions [6]. Data warehouse architectures often incorporate specialized data marts tailored to specific domains or diseases by integrating clinical, geographic, and environmental data [7]. For instance, in dengue-endemic regions like Bangladesh, where seasonal outbreaks pose significant healthcare challenges, a dedicated dengue-focused data mart can enhance the opportunity to detect outbreaks, plan resource allocation, and conduct epidemiological research [8, 9]. Data mining techniques such as clustering and association can also uncover hidden patterns in health data, aiding outbreak detection [10, 11].
The choice of SQL (e.g., PostgreSQL) vs. NoSQL (e.g., HBase) is also critical for managing healthcare data. SQL ensures ACID compliance, while NoSQL offers better scalability for semi-structured data. A recent study in Bangladesh suggests hybrid approaches improve performance for large-scale health analytics [12].
@@
In earlier work [13], existing national health data management systems have been analyzed, along with the design and limitations of Bangladesh’s NCDW, and the need to strengthen its performance, security, and integration of heterogeneous data has been explored. Our previous work [14, 15] proposed a privacy-preserving NCDW architecture that leveraged the key-based algorithm to anonymize patient data while enabling secure linkages across fragmented sources. Although this solution addressed fundamental privacy concerns, limitations were identified in areas such as scalability, interoperability, and the incorporation of external datasets for comprehensive analysis. To overcome these challenges, this study presents an enhanced NCDW architecture with the following key improvements:
@@ II Proposed Architecture
The proposed National Clinical Data Warehouse (NCDW) architecture, shown in Fig. 1, can be divided into five stages: (1) Data Integration, (2) Data Standardization, (3) Analytical Infrastructure, (4) Centralized Data Management, and (5) Analytical Dashboard.
Figure 1: Wrapper-based Architecture of NCDW Supporting Dengue Data Mart
@@ II-C Centralized Data Warehousing and Management
TABLE I: Details of newly added fact and dimension tables
@@
Cube Hierarchy and Aggregation: As illustrated in Fig. 2, aggregated data from the schema enables the creation of multi-dimensional cubes for diagnosis and disease analysis. The 5D Clinical Test Data Cube, depicted in Fig. 3, supports national CDW development, with a cube hierarchy ranging from the 0-D apex cube, summarizing national health statistics, to the 4-D base cube for granular data. The system precomputes common analytics, such as patient counts by diagnosis, procedure tracking, and retesting statistics, ensuring fast and efficient query responses. These preaggregated datasets enhance knowledge discovery, prediction, and dataset generation, aligning with the warehouse’s goal of enabling real-time insights.
Figure 2: CUBE Hierarchy In Disease Analysis
Figure 3: Clinical Test Data Cube
Figure 4: Star Schema of NCDW Supporting Dengue Data Mart
@@ II-F Name Matching for Data Consistency
To address patient identification challenges, we proposed a name-matching algorithm [1] that generates phonetic identifiers to handle spelling variations and pronunciation differences. It converts names into a fixed-length four-character code, reducing ambiguity and ensuring standardized comparison. For instance, “Sobuj Chowdhury” and “Sabuj Chaudhury” produce identical Soundex codes, enabling accurate record linkage. This approach minimizes data inconsistencies and improves patient record management in resource-constrained healthcare systems, as shown in Table IV.
@@ III-A1 Load Estimation and Storage Requirements
$\text{Size}=R\times d\times S_{r}$ (2)
@@
TABLE VII: Estimated size of NCDW for different time spans
@@
TABLE VIII: Execution Times (in minutes) for Cube Sizes 3 and 4
@@ III-B3 Outbreak Detection
Figure 11: Distribution of the most frequently conducted diagnostic tests for dengue-positive patients, illustrating the high reliance on CBC, Dengue NS1, Electrolyte, and creatinine serum tests for disease monitoring and management.
Fig. 11 presents the top four most frequently conducted laboratory tests for dengue-positive cases. CBC, Dengue NS1, and electrolyte serum tests dominate, reflecting their critical role in dengue diagnosis and patient monitoring.

### v2

@@ Abstract
Index Terms— Dengue, Datamart, Health Informatics, Data Mining, Warehouse, Electronic Health Records
@@ I Introduction & Related Works
A National Clinical Data Warehouse (NCDW) offers a viable approach to consolidating multi-institutional data into a privacy-preserving environment. By unifying patient information, laboratory results, and clinical observations, an NCDW can enhance large-scale health analytics and support more informed policy decisions [6]. Data warehouse architectures often incorporate specialized data marts tailored to specific domains or diseases by integrating clinical, geographic, and environmental data [7]. For instance, in dengue-endemic regions like Bangladesh, where seasonal outbreaks pose significant healthcare challenges, a dedicated dengue-focused data mart can enhance the opportunity to detect outbreaks, plan resource allocation, and conduct epidemiological research [8, 9]. Data mining techniques such as clustering and association can also uncover hidden patterns in health data, aiding outbreak detection [10, 11, 12].
The choice of SQL (e.g., PostgreSQL) vs. NoSQL (e.g., HBase) is also critical for managing healthcare data. SQL ensures ACID (atomicity, consistency, isolation, durability) compliance, while NoSQL offers better scalability for semi-structured data. A recent study in Bangladesh suggests hybrid approaches improve performance for large-scale health analytics [13].
@@
In earlier work [14], existing national health data management systems have been analyzed, along with the design and limitations of Bangladesh’s NCDW, and the need to strengthen its performance, security, and integration of heterogeneous data has been explored. Our previous works [15, 16] proposed a privacy-preserving NCDW architecture that leveraged the key-based algorithm to anonymize patient data while enabling secure linkages across fragmented sources. Although this solution addressed fundamental privacy concerns, limitations were identified in areas such as scalability, interoperability, and the incorporation of external datasets for comprehensive analysis. To overcome these challenges, this study presents an enhanced NCDW architecture with the following key improvements:
@@ II Proposed Architecture
The proposed National Clinical Data Warehouse (NCDW) architecture, shown in Fig. 1, can be divided into four stages: (1) Data Integration, (2) Data Standardization, (3) Centralized Data Management, and (4) Analytical Infrastructure and Dashboard.
Figure 1: Wrapper-based architecture of NCDW supporting disease-specific data mart.
@@ II-C Centralized Data Warehousing and Management
TABLE I: Details of newly added fact and dimension tables.
@@
Cube Hierarchy and Aggregation: As illustrated in Fig. 3, aggregated data from the schema enables the creation of multi-dimensional cubes for diagnosis and disease analysis. The 5D clinical test data cube shown in Fig. 4 supports NCDW development, with a cube hierarchy ranging from the 0-D apex cube, summarizing national health statistics, to the 4-D base cube for granular data. The system precomputes common analytics, such as patient counts by diagnosis, procedure tracking, and retesting statistics, ensuring fast and efficient query responses. These preaggregated datasets enhance knowledge discovery, prediction, and dataset generation, aligning with the warehouse’s goal of enabling real-time insights.
Figure 2: Star schema of NCDW supporting dengue data mart.
Figure 3: CUBE hierarchy in disease analysis.
Figure 4: Clinical test data cube.
@@ II-F Name Matching for Data Consistency
To address patient identification challenges, we proposed a name-matching algorithm, Soundex [1] that generates phonetic identifiers to handle spelling variations and pronunciation differences. It converts names into a fixed-length four-character code, reducing ambiguity and ensuring standardized comparison. For instance, “Sobuj Chowdhury” and “Sabuj Chaudhury” produce identical Soundex codes (S132 C360), enabling accurate record linkage. This approach minimizes data inconsistencies and improves patient record management in resource-constrained healthcare systems, as shown in Table IV.
@@ III-A1 Load Estimation and Storage Requirements
$\vspace{-2mm}\text{Size}=R\times d\times S_{r}$ (2)
@@
TABLE VII: Estimated size of NCDW for different time spans.
@@
TABLE VIII: Execution times (in minutes) for cube sizes 3 and 4.